# Deep Class-Incremental Learning: A Survey

### v1

@@ 5 Future Directions
CIL with Complex Inputs: In the real world, data is often with complex format [202, 203], e.g., few-shot [204], imbalanced [136], with noisy label [205], weak supervised [206], multi modal [207], concept drift [13], novel classes [208, 209] etc. Incremental learning methods should be able to handle these real-world scenarios for better generalizability. Figure 13 depicts possible future applications of class-incremental learning.
@@
Specifically, [104, 106, 210, 211] address the few-shot class-incremental learning problem, where the model is required to be adapted to incoming few-shot classes. [212, 213, 214] propose to tackle the long-tailed class-incremental learning problem, where the head classes are easy to collect with adequate instances while tail classes are scarce. Recently, with the prosperity of CLIP [147], incrementally training the language vision models to handle multi-modal data streams [188, 189] is becoming popular. Since the labeling cost is always expensive in the real world, there are some works addressing training CIL models in a semi-supervised [215] or unsupervised manner [171]. [33] proposes a unified framework to handle CIL with concept drift. If the test dataset contains instances from unknown classes, open-set recognition [208] and novel class discovery [216, 217] can equip the model with the detection ability, which refers to the open-world recognition problem [218]. Besides, test-time adaptation [219] enables the model to handle these drifts on-the-fly. Lastly, real-world data may emerge hierarchically, and the labels could evolve from coarse-grained to fine-grained [220]. Learning incremental models with refined concepts [221, 222, 223, 224, 225] is also vital to building real-world learning systems.
@@
CIL with General Data Stream: Current CIL methods have a set of restrictions on the data stream, e.g., saving exemplars for rehearsal, undergoing multi-epoch offline training within an incremental task, etc. Future class-incremental learning algorithms should be able to conduct fully online training [226, 227, 228, 229, 230] without requiring the task boundaries [231, 232, 233]. On the other hand, saving exemplars from the history may violate user privacy in some cases. To this end, exemplar-free CIL [58, 139, 138, 57] should be conducted to enable the model to be adapted without the help of exemplars. Lastly, most CIL methods rely on the number of base classes to define hyper-parameters in model optimization, where more base classes require larger stability and fewer base classes require larger plasticity. Therefore, designing the algorithm to handle CIL problems given any base classes [234] is also important to the real world.
@@
CIL with Any Memory Budgets: Dynamic networks have obtained impressive performance in recent years [21, 20]. However, most of them require an extra memory budget to save the external backbones. In the real world, a good CIL algorithm should handle different budget restrictions. For example, an algorithm should be able to learn with high-performance computers or with edge devices (e.g., smartphones), and both scenarios are essential. Hence, deploying and comparing CIL models in the real world should take the memory budgets into consideration. The performance measures of AUC-A/L [22] are proper solutions to holistically compare different methods given any memory budgets. On the other hand, future CIL methods are encouraged to handle specific learning scenarios. For example, [235] addresses training CIL systems under resource-limited scenarios.
@@
CIL with Pre-trained Models: Recently, pre-trained models have shown to work competitively with their strong transferability, especially for ViT-based methods [73, 72, 74, 75]. The pre-trained models provide generalizable features for the downstream tasks, enabling the incremental model to adjust with minimal cost [145]. CIL with pre-trained models is a proper way to handle real-world incremental applications from an excellent starting point. However, since the final target for incremental learning is to build a generalizable feature continually [236], some will argue that pre-trained models weaken the difficulty of incremental learning. From this perspective, designing proper algorithms to train the CIL model from scratch is more challenging. On the other hand, pre-trained language-vision models [147, 237] have shown powerful generalizability in recent years, and exploring the ensemble of various pre-trained models [238, 239] is also an exciting topic.
@@
Analyzing the Reason Behind Catastrophic Forgetting: Model rectification-based methods aim to reduce the inductive bias in the CIL model. Recently, more works have tried to analyze the reason for forgetting in CIL. [172, 246] find that batch normalization layers are biased when sequentially trained, which triggers the different activation of old and new classes. [138] finds that representations with large eigenvalues transfer better and suffer less forgetting, and [226] finds that eigenvalues can be enlarged by maximizing the mutual information between old and new features. [247] theoretically decomposes the CIL problem into within-task and task-id predictions. It further proves that good within-task and task-id predictions are necessary and sufficient for good CIL performances. [248] finds that applying data replay causes the newly added classes’ representations to overlap significantly with the previous classes, leading to highly disruptive parameter updates. [249] empirically finds that the amount of forgetting correlates with the geometrical properties of the convergent points. It would be interesting to explore more reasons for catastrophic forgetting theoretically and empirically in the future.
@@ 6 Conclusion
Real-world applications are often faced with streaming data, with which the model should be incrementally updated without catastrophic forgetting. In this paper, we provide a comprehensive survey about deep class-incremental learning by categorizing them into three categories and six subcategories taxonomically and chronologically. Additionally, we provide a holistic comparison among different methods on several publicly available datasets. With these results, we discuss the insights and summarize the common rules to inspire future research. Finally, we highlight an important factor in CIL comparison, namely memory budget, and advocate evaluating different methods holistically by emphasizing the effect of memory budgets. We provide a comprehensive evaluation of different methods given specific budgets as well as some new performance measures. We expect this survey to provide an effective way to understand current state-of-the-art and speed up the development of the class-incremental learning field.
@@ A.1 Detailed Results of CIFAR100
As we can infer from these figures, L2P consistently obtains the best performance among all settings. However, it must be noted that L2P relies on the ImageNet-21K pre-trained Vision Transformer with 85M parameters. At the same time, typical methods train ResNet32 with 0.46M parameters from scratch. Hence, comparing pre-trained model based methods to typical methods is unfair. We also notice that dynamic networks, e.g., DER and MEMO obtain better performance at the cost of more parameters. Therefore, we report Table V to point out that parameter cost is sometimes ignored in the benchmark comparison, which shall lead to unfair results. These comparison details also motivate us to design fair comparison protocols for class-incremental learning.
@@ A.2 Detailed Results of ImageNet100
Figure 15: Incremental top-1 accuracy of different methods on ImageNet100. Legends are shown in (f).
@@ A.3 ImageNet Top-5 Results
Figure 16: Incremental top-5 accuracy of different methods on ImageNet100. Legends are shown in (d).
@@
Figure 17: Incremental top-5 accuracy of different methods on ImageNet1000. Legends are shown in (b).
@@
Appendix B Supplied Visualization Results
In this section, we report the full visualization results in the CIFAR100 comparison. Specifically, we first show the confusion matrix of 15 methods in Figure 18. As we can infer from these figures, finetune and EWC forget most knowledge from former classes, while LwF shows less forgetting with knowledge distillation. When comparing LwF to EWC, we can find that knowledge distillation is a more suitable solution to resist catastrophic forgetting than parameter regularization without saving exemplars. When equipping the model with exemplars, the diagonal elements become brighter, implying that forgetting is alleviated. The performance of Replay shows more steady improvement than finetune, indicating the critical role of exemplars in CIL. However, when comparing iCaRL to GEM, we can find that data regularization shows inferior performance than knowledge distillation in the efficiency of using exemplars. Moreover, BiC and WA improve the performance of iCaRL by rectifying the bias in the model, indicating that these methods can be orthogonally equipped to other CIL methods for rectification. Lastly, dynamic networks (i.e., DER, FOSTER, MEMO, and DyTox) still show the most competitive performance in these visualizations, indicating that the former backbones help the model retrieve old features and resist forgetting.
Additionally, we visualize the weight norm of the fully-connected layers in Figure 19 to show the bias introduced by incremental learning. As we can infer from these figures, when sequentially finetuning the model, the weight norm of new classes will dominate the learning process due to data imbalance. As shown in Figure 19(a), the weight norm of finetune has a stair-step shape. Since the features are passed through the ReLU layer to avoid negative ones, the larger weight norm of new classes will result in an imbalance in the prediction results. These results are consistent with the confusion matrix in Figure 18, and we can find that most CIL methods can relieve the imbalance in the fully connected layers.
@@ Appendix B Supplied Visualization Results
(a) Average accuracy curve of CIFAR100 Base0 Inc10.
(b) Average accuracy curve of ImageNet100 Base50 Inc5.
(c) Last accuracy curve of CIFAR100 Base0 Inc10.
(d) Last accuracy curve of ImageNet100 Base50 Inc5.
@@
Appendix C Supplied Results in Memory-Aligned Comparison
In the main paper, we report the incremental performance curve with aligned memory to DER. In this section, we report the detailed setting of those figures, including the number of exemplars, the number of parameters, the total memory size, and the detailed performance in Table VII. Specifically, since DER and MEMO require an extra model budget than other methods, we equip more exemplars to other methods to make sure the total memory size is aligned to the same scale.
@@ D.1 Necessity of AUC in CIL
We report the full performance-memory curve in Figure 20, including the average accuracy-memory curve and last accuracy-memory curve. In each figure, the X-coordinate indicates the memory budget of the specific model, and the Y-coordinate represents the corresponding performance of the model. To highlight the importance of the AUC measure, we begin with a simple question:
@@ D.2 Implementation Details
Discussion about selected methods: In this part, we choose ten methods for the comparison, i.e., Finetune, EWC, LwF, RMM, DyTox, and L2P are not included in the evaluation. Firstly, non-exemplar-based methods (Finetune, EWC, LwF) are unsuitable for the evaluation since the extendability relies on adding extra exemplars. Secondly, RMM advocates another evaluation protocol in terms of the memory, which is incompatible with ours. Lastly, DyTox and L2P rely on the vision transformer as the backbone, which have a much larger memory scale than benchmark backbones. Directly comparing ResNet to ViT may be unfair since these backbones have different characteristics in model optimization.
@@ D.2.2 ImageNet100 Implementations
ImageNet100 with 329 MB Memory Size: The implementations are shown in Figure 26. $329$ MB is a relatively small memory size. Since we need to align the total budget of these methods, we are only able to use small backbones for DER and MEMO. These small backbones, i.e., ConvNet with four convolutional layers, have much fewer parameters than ResNet18 and saving 10 ConvNets matches the memory size of a single ResNet18. We can infer from the figure that DER and MEMO are restricted by the learning ability of the inferior backbones, which perform poorly in the base task.

### v2

@@ 5 Future Directions
CIL with Complex Inputs: In the real world, data is often with complex format, e.g., few-shot [27], imbalanced [28], weak supervised [184, 214], multi modal [180, 183], concept drift [10], novel classes [215] etc. CIL methods should be able to handle these real-world scenarios for better generalizability. Specifically, [27, 104, 121, 115, 216] address the few-shot CIL problem, where the model is required to be adapted to incoming few-shot classes. [141, 28] propose to tackle the long-tailed CIL problem, where the head classes are easy to collect with adequate instances while tail classes are scarce. Recently, with the prosperity of CLIP [180], incrementally training the language vision models to handle multi-modal data streams [217] is becoming popular. Since the labeling cost is always expensive in the real world, there are some works addressing training CIL models in a semi-supervised or unsupervised manner [184]. [148] proposes a unified framework to handle CIL with concept drift. If the test dataset contains instances from unknown classes, open-set recognition [215] and novel class discovery [218] can equip the model with the detection ability, which refers to the open-world recognition problem [219]. Lastly, real-world data may emerge hierarchically, and the labels could evolve from coarse-grained to fine-grained. CIL with refined concepts [220] is also vital to building real-world learning systems.
@@
CIL with General Data Stream: Current CIL methods have a set of restrictions on the data stream, e.g., saving exemplars for rehearsal, undergoing multi-epoch offline training within an incremental task, etc. Future CIL algorithms should be able to conduct fully online training [221] without requiring the task boundaries [222]. On the other hand, saving exemplars from the history may violate user privacy in some cases. To this end, exemplar-free CIL [50, 150, 223, 224, 225] should be conducted to enable the model to be adapted without the help of exemplars. Lastly, most CIL methods rely on the number of base classes to define hyper-parameters in model optimization, where more base classes require larger stability and fewer base classes require larger plasticity. Therefore, designing the algorithm to handle CIL problems given any base classes [226] is also essential to the real world.
@@
CIL with Any Memory/Computational Budgets: Dynamic networks have obtained impressive performance in recent years [16]. However, most of them require an extra memory budget to save the external backbones. In the real world, a good CIL algorithm should handle different budget restrictions. For example, an algorithm should be able to learn with high-performance computers or with edge devices (e.g., smartphones), and both scenarios are essential. Hence, deploying and comparing CIL models in the real world should take the memory budgets into consideration. The performance measures of AUC-A/L [17] are proper solutions to holistically compare different methods given any memory budgets. On the other hand, future CIL methods are also encouraged to handle specific learning scenarios, e.g., [227] addresses training CIL systems under resource-limited scenarios. Another important characteristic is the computational budget [228]. For realistic scenarios with high-throughput streams, computational bottlenecks impose implicit constraints on learning from past samples that can be too many to be revisited during training. In that case, we can only update the model with limited iterations and cannot tackle all the data. Hence, designing computational-efficient algorithms in real-world applications remains a promising direction for relating CIL to realistic scenarios.
@@
CIL with Pre-trained Models: Recently, pre-trained models have shown to work competitively with their strong transferability, especially for ViT-based methods [70, 69, 71, 72]. The pre-trained models provide generalizable features for the downstream tasks, enabling the incremental model to adjust with minimal cost [179]. CIL with pre-trained models is a proper way to handle real-world incremental applications from an excellent starting point. However, since the final target for incremental learning is to build a generalizable feature continually, some will argue that pre-trained models weaken the difficulty of incremental learning. From this perspective, designing proper algorithms to train the CIL model from scratch is more challenging. On the other hand, pre-trained language-vision models [180] have shown powerful generalizability in recent years, and exploring the ensemble of various pre-trained models is also an exciting topic.
@@
Analyzing the Reason Behind Catastrophic Forgetting: Model rectification-based methods aim to reduce the inductive bias in the CIL model. Recently, more works have tried to analyze the reason for forgetting in CIL. [196] suggests that batch normalization layers are biased when sequentially trained, which triggers the different activation of old and new classes. [223] finds that representations with large eigenvalues transfer better and suffer less forgetting, and [221] shows that eigenvalues can be enlarged by maximizing the mutual information between old and new features. [235] theoretically decomposes the CIL problem into within-task and task-id predictions. It further proves that good within-task and task-id predictions are necessary and sufficient for good CIL performances. [236] finds that applying data replay causes the newly added classes’ representations to overlap significantly with the previous classes, leading to highly disruptive parameter updates. [200] empirically shows that the amount of forgetting correlates with the geometrical properties of the convergent points. It would be interesting to explore more reasons for catastrophic forgetting theoretically and empirically in the future.
@@ 6 Conclusion
Real-world applications often face streaming data, with which the model should be incrementally updated without catastrophic forgetting. In this paper, we provide a comprehensive survey about class-incremental learning by dividing them into seven categories taxonomically and chronologically. Additionally, we provide a holistic comparison among different methods on several publicly available datasets. With these results, we discuss the insights and summarize the common rules to inspire future research. Finally, we highlight an important factor in CIL comparison, namely memory budget, and advocate evaluating different methods holistically by emphasizing the effect of memory budgets. We provide a comprehensive evaluation of different methods given specific budgets as well as some new performance measures. We expect this survey to provide an effective way to understand current state-of-the-art and speed up the development of the CIL field.
Supplementary Material
The supplementary material mainly contains additional experiments that cannot be reported due to page limit, which is organized as follows:
  • Section A provides extra results in benchmark comparison, including detailed results on CIFAR100, ImageNet100, and Top-5 accuracy on ImageNet.
  • Section B provides visualization results on the confusion matrix and weight norm.
  • Section C provides extra results on other settings in the memory-aligned comparison.
  • Section D reports the detailed implementations of the AUC-A/L measure, including the per-point setting of memory, backbone, and incremental performance.
  • Section E reports the incremental learning performance (per-stage accuracy) of different methods in the benchmark comparison.
  • Section F reports the extra evaluation measures, i.e., forgetting and intransigence among all compared methods.
  • Section G conducts experiments on the type of replay strategies, including direct replay, feature replay, and generative replay.
  • Section H gives a detailed introduction about selected compared methods in the main paper.
  • Section I discusses the taxonomy in the main paper and introduces the related topics about incremental object detection and semantic segmentation.
  • Section J explores further analysis in class-incremental learning, including the influence of exemplars and computational efficiency.
(a) CIFAR100 Base0 Inc5
(b) CIFAR100 Base0 Inc10
(c) CIFAR100 Base0 Inc20
(d) CIFAR100 Base50 Inc50
(e) CIFAR100 Base50 Inc10
(f) CIFAR100 Base50 Inc25
Figure 10: Incremental accuracy of different methods on CIFAR100.
@@ A.1 Detailed Results of CIFAR100
As we can infer from these figures, L2P consistently obtains the best performance among all settings. However, it must be noted that L2P relies on the ImageNet-21K pre-trained Vision Transformer with 85M parameters. At the same time, typical methods train ResNet32 with 0.46M parameters from scratch. Hence, comparing pre-trained model-based methods to typical methods is unfair. We also notice that dynamic networks, e.g., DER and MEMO obtain better performance at the cost of more parameters. Therefore, we report Table V to point out that parameter cost is sometimes ignored in the benchmark comparison, which shall lead to unfair results. These comparison details also motivate us to design fair comparison protocols for class-incremental learning.
@@ A.2 Detailed Results of ImageNet100
Figure 11: Incremental top-1 accuracy of different methods on ImageNet100.
@@ A.3 ImageNet Top-5 Results
Figure 12: Incremental top-5 accuracy of different methods on ImageNet100.
@@
Figure 13: Incremental top-5 accuracy of different methods on ImageNet1000.
@@
Appendix B Visualization Results
In this section, we report the full visualization results in the CIFAR100 comparison. Specifically, we first show the confusion matrix of 15 methods in Figure 14. As we can infer from these figures, Finetune and EWC forget most knowledge from former classes, while LwF shows less forgetting with knowledge distillation. When comparing LwF to EWC, we find that knowledge distillation is a more suitable solution to resist catastrophic forgetting than parameter regularization without saving exemplars. When equipping the model with exemplars, the diagonal elements become brighter, implying that forgetting is alleviated. The performance of Replay shows more steady improvement than finetune, indicating the critical role of exemplars in CIL. However, when comparing iCaRL to GEM, we can find that data regularization shows inferior performance than knowledge distillation in the efficiency of using exemplars. Moreover, BiC and WA improve the performance of iCaRL by rectifying the bias in the model, indicating that these methods can be orthogonally equipped to other CIL methods for rectification. Lastly, dynamic networks (i.e., DER, FOSTER, MEMO, and DyTox) still show the most competitive performance in these visualizations, indicating that the former backbones help the model retrieve old features and resist forgetting.
Additionally, we visualize the weight norm of the fully-connected layers in Figure 15 to show the bias introduced by incremental learning. As we can infer from these figures, when sequentially finetuning the model, the weight norm of new classes will dominate the learning process due to data imbalance. As shown in Figure 15(a), the weight norm of finetune has a stair-step shape. Since the features are passed through the ReLU layer to avoid negative ones, the larger weight norm of new classes will result in an imbalance in the prediction results. These results are consistent with the confusion matrix in Figure 14, and we find that most CIL methods can relieve the imbalance in the fully connected layers.
@@ Appendix B Visualization Results
Appendix C Supplied Results in Memory-Aligned Comparison
In the main paper, we report the incremental performance curve with aligned memory to DER. In this section, we report the detailed setting of those figures, including the number of exemplars, the number of parameters, the total memory size, and the detailed performance in Table VII. Specifically, since DER and MEMO require an extra model budget compared to other methods, we equip more exemplars to other methods to make sure the total memory size is aligned to the same scale.
(a) Average accuracy curve of CIFAR100 Base0 Inc10
(b) Average accuracy curve of ImageNet100 Base50 Inc5
(c) Last accuracy curve of CIFAR100 Base0 Inc10
(d) Last accuracy curve of ImageNet100 Base50 Inc5
@@ D.1 Necessity of AUC in CIL
We report the full performance-memory curve in Figure 16, including the average and last accuracy-memory curves. In each figure, the X-coordinate indicates the memory budget of the specific model, and the Y-coordinate represents the corresponding performance of the model. To highlight the importance of the AUC measure, we begin with a simple question:
@@ D.2 Implementation Details
Discussion about selected methods: In this part, we choose ten methods for the comparison, i.e., Finetune, EWC, LwF, RMM, DyTox, and L2P are not included in the evaluation. Firstly, non-exemplar-based methods (Finetune, EWC, LwF) are unsuitable for the evaluation since the extendability relies on adding extra exemplars. Secondly, RMM advocates another evaluation protocol in terms of the memory, which is incompatible with ours. Lastly, DyTox and L2P rely on the vision transformer as the backbone, which has a much larger memory scale than benchmark backbones. Directly comparing ResNet to ViT may be unfair since these backbones have different characteristics in model optimization.
@@ D.2.2 ImageNet100 Implementations
ImageNet100 with 329 MB Memory Size: The implementations are shown in Figure 22. $329$ MB is a relatively small memory size. Since we need to align the total budget of these methods, we are only able to use small backbones for DER and MEMO. These small backbones, i.e., ConvNet with four convolutional layers, have much fewer parameters than ResNet18, and saving 10 ConvNets matches the memory size of a single ResNet18. We can infer from the figure that DER and MEMO are restricted by the learning ability of the inferior backbones, which perform poorly in the base task.